# Efficient Estimation of the Central Mean Subspace via Smoothed Gradient Outer Products

### v1

@@ Abstract
We consider the problem of sufficient dimension reduction (SDR) for multi-index models. The estimators of the central mean subspace in prior works either have slow (non-parametric) convergence rates, or rely on stringent distributional conditions (e.g., the covariate distribution $P_{\boldsymbol{X}}$ being elliptical symmetric). In this paper, we show that a fast parametric convergence rate of form $C_{d}\cdot n^{-1/2}$ is achievable via estimating the expected smoothed gradient outer product, for a general class of distribution $P_{\boldsymbol{X}}$ admitting Gaussian or heavier distributions. When the link function is a polynomial with a degree of at most $r$ and $P_{\boldsymbol{X}}$ is the standard Gaussian, we show that the prefactor depends on the ambient dimension $d$ as $C_{d}\propto d^{r}$..
@@ 1 Introduction
Our result relies on estimating outer products of a smoothed version of $\nabla f$, where the smoothing itself—induced by careful reweighting of the training sample—leverages a first-order Stein’s Lemma to guarantee full recovery of $\operatorname{col}(\boldsymbol{U})$. Similar to other existing analyses, we use knowledge of $P_{\boldsymbol{X}}$ to properly reweighting of the training sample.
@@ 1.1 Prior Theoretical Works
As the name suggests, the class of inverse regression methods leverage functionals of the inverse relation $\boldsymbol{X}\mid Y$ that may reveal the index space or a subspace thereof. Although these methods are initially designed to estimate the central subspace, they can also serve the purpose of estimating the central mean subspace when the covariates $\boldsymbol{X}$ and the error $\varepsilon=Y-g(\boldsymbol{X})$ are independent, under which case these two subspaces are identical. For instance, Sliced Inverse Regression (SIR) by [21] estimates $\boldsymbol{M}=\text{Cov}[\operatorname*{\mathbb{E}}[\boldsymbol{X}\mid Y]]$ and Sliced Average Variance Estimator (SAVE) by [9] estimates $\boldsymbol{M}=\operatorname*{\mathbb{E}}[(\boldsymbol{I}_{d}-\text{Cov}[\boldsymbol{X}\mid Y])^{2}]$. By carefully estimating means or covariances over slices, i.e., partitions of the $Y$ space, parametric convergence rates of the form $O(n^{-1/2})$ may be guaranteed, albeit under restrictive distributional conditions. For instance, SIR imposes a so-called linear conditional mean condition (where $\operatorname*{\mathbb{E}}[\boldsymbol{X}\mid\boldsymbol{U}^{{\scriptscriptstyle{\mathsf{T}}}}\boldsymbol{X}]$ is linear in $\boldsymbol{U}^{{\scriptscriptstyle{\mathsf{T}}}}\boldsymbol{X}$) in order to ensure $\operatorname{col}(\boldsymbol{M})\subset\operatorname{col}(\boldsymbol{U})$, i.e., in order to guarantee partial recovery of $\operatorname{col}(\boldsymbol{U})$; if in addition we impose a so-called random conditional projection condition (requiring for all $v\in\mathbb{R}^{d}$ that $\operatorname*{\mathbb{E}}[\boldsymbol{v}^{{\scriptscriptstyle{\mathsf{T}}}}\boldsymbol{X}\mid Y]$ is not constant as a r.v.) then one can ensure $\operatorname{col}(\boldsymbol{M})=\operatorname{col}(\boldsymbol{U})$. These distributional conditions, by restricting the relation between $P_{\boldsymbol{X}}$ and the index space $\operatorname{col}(\boldsymbol{U})$, rule out some rather benign settings: for instance, for $k=1$, it is well known that under Gaussian $P_{\boldsymbol{X}}$, these conditions cannot hold for even link functions such as $f(z)=z^{2}$ (also referred as phase retrieval problem). Similar examples of failure modes are discussed for SAVE in Appendix A. Additional discussions and examples for non-exhaustiveness issues can be found in [14] and [19].
@@ 1.2 Detailed Contributions
Our main result states that, for a general class of distributions $P_{\boldsymbol{X}}$ admitting Gaussians and heavier tail distributions, our estimator converges at a parametric rate of the form $C_{d}\cdot n^{-1/2}$. Finally, to elucidate the leading constant $C_{d}$, we consider the special case of polynomial links $g$—which has received much recent interest—and show that it behaves as $d^{O(r)}$ for polynomials of degree at most $r$ (Corollary 3.19), in line with recent lower-bounds of [12] for certain classes of computationally tractable procedures.
@@ 2.2 Central Mean Subspace
$g(\boldsymbol{X})=f(\boldsymbol{U}^{{\scriptscriptstyle{\mathsf{T}}}}\boldsymbol{X})$ (1)
@@
The distance with optimal rotation between $\boldsymbol{U}$ and its estimates ^$\hat{}\boldsymbol{U}$ is defined as:
$d(\hat{}\boldsymbol{U},\boldsymbol{U}):=\min_{\boldsymbol{R}\in\mathcal{O}^{k\times k}}\lVert\hat{}\boldsymbol{U}\boldsymbol{R}-\boldsymbol{U}\rVert_{\operatorname{op}}\,,$
@@
$c\cdot d(\widehat{\boldsymbol{U}},\boldsymbol{U})\leq\sin_{\Theta}(\widehat{\boldsymbol{U}},\boldsymbol{U})\leq C\cdot d(\widehat{\boldsymbol{U}},\boldsymbol{U})_{\operatorname{op}}.$
@@ 3 Main Results
We first introduce a surrogate object $\overline{\boldsymbol{M}}$ in Section 3.1 below, and show that it indeed is exhaustive, i.e., it recovers all of $\operatorname{col}({\bf U})$. Following this, we derive a separate surrogate object $\widetilde{\boldsymbol{M}}$ in Section 3.3 which approximates $\overline{\boldsymbol{M}}$ and turns out to be simpler to estimate and retains the same exhaustiveness properties. Our main algorithm then estimates $\widetilde{\boldsymbol{M}}$, and returns the column space $\hat{U}$ of this estimate. Our main result of Theorem 3.13 which bounds the estimation error is presented in Section 3.4, followed by an instantiation for polynomial link functions in Section 3.5.
@@ 3.1 Expected Smoothed Gradient Outer Product
$\overline{\boldsymbol{M}}:=\operatorname*{\mathbb{E}}_{\boldsymbol{\theta}\sim\mathcal{N}(\boldsymbol{0}_{d},\sigma_{\theta}^{2}\boldsymbol{I}_{d})}\left[\boldsymbol{\beta}_{h}(\boldsymbol{\theta})\boldsymbol{\beta}_{h}(\boldsymbol{\theta})^{{\scriptscriptstyle{\mathsf{T}}}}\right]=\boldsymbol{U}\left(\operatorname*{\mathbb{E}}_{\boldsymbol{\theta}\sim\mathcal{N}(\boldsymbol{0}_{d},\sigma_{\theta}^{2}\boldsymbol{I}_{d})}\left[\overline{\nabla}_{h}f(\boldsymbol{\theta})\overline{\nabla}_{h}f(\boldsymbol{\theta})^{\scriptscriptstyle{\mathsf{T}}}\right]\right)\boldsymbol{U}^{\scriptscriptstyle{\mathsf{T}}}.$ (2)
@@ 3.2 Average Smoothed Gradient Outer Product
$\widetilde{\boldsymbol{M}}:=\frac{1}{m}\sum_{j=1}^{m}\boldsymbol{\beta}_{h}(\boldsymbol{\theta}_{j})\boldsymbol{\beta}_{h}(\boldsymbol{\theta}_{j})^{{\scriptscriptstyle{\mathsf{T}}}}=\boldsymbol{U}\left(\frac{1}{m}\sum_{j=1}^{m}\left[\overline{\nabla}_{h}f(\boldsymbol{\theta}_{j})\overline{\nabla}_{h}f(\boldsymbol{\theta}_{j})^{\scriptscriptstyle{\mathsf{T}}}\right]\right)\boldsymbol{U}^{\scriptscriptstyle{\mathsf{T}}}.$ (3)
@@
The smooth gradient ¯$\overline{\nabla}_{h}f$ satisfies:
@@ 3.3 Estimation of CMS
For intuition, notice that we have $\boldsymbol{\beta}_{h}(\boldsymbol{\theta})=\operatorname*{\mathbb{E}}_{\boldsymbol{Z}\sim\mathcal{N}(\boldsymbol{0},h^{2}\boldsymbol{I}_{d})}\nabla g(\boldsymbol{Z}+\boldsymbol{\theta})=\operatorname*{\mathbb{E}}_{\boldsymbol{Z}\sim\mathcal{N}(\boldsymbol{\theta},h^{2}\boldsymbol{I}_{d})}\nabla g(\boldsymbol{Z})$. Thus, suppose for a moment we were to sample $\boldsymbol{X}\sim\mathcal{N}(\boldsymbol{\theta},h^{2}\boldsymbol{I}_{d})$, then we would have
@@
$\rho_{h}(\boldsymbol{x};{\boldsymbol{\theta}}):=\frac{\varphi_{h}(\boldsymbol{x}-\boldsymbol{\theta})}{p(\boldsymbol{x})}\,,\quad\text{where }\varphi_{h}(\cdot)\text{ denotes the density of $\mathcal{N}(\boldsymbol{0}_{d},h^{2}\boldsymbol{I}_{d})$ }.$ (4)
@@
$\hat{\boldsymbol{\beta}}_{h}(\boldsymbol{\theta}_{j};\mathcal{D}_{j,\ell}):=\frac{h^{-2}}{|\mathcal{D}_{j,\ell}|}\sum_{(\boldsymbol{X},Y)\in\mathcal{D}_{j,\ell}}\rho_{h}(\boldsymbol{X};\boldsymbol{\theta}_{j})\cdot Y\cdot(\boldsymbol{X}-\boldsymbol{\theta}_{j}).$ (5)
@@
Notice that, by the independence of the two estimates of $\boldsymbol{\beta}_{h}(\boldsymbol{\theta}_{j})$, we also immediately have that $\hat{\boldsymbol{\beta}}_{j,1}\hat{\boldsymbol{\beta}}_{j,2}^{{\scriptscriptstyle{\mathsf{T}}}}={\boldsymbol{\beta}}_{h}(\boldsymbol{\theta}_{j}){\boldsymbol{\beta}}_{h}(\boldsymbol{\theta}_{j})^{{\scriptscriptstyle{\mathsf{T}}}}$. This leads to the estimator of $\widetilde{\boldsymbol{M}}$ given below:
@@ 3.4 Upper-bound on Estimation Error
The noise has a subgaussian tail, i.e., $\exists C,\sigma_{Y}^{2}>0$ such that $\forall\boldsymbol{x}\in\mathbb{R}^{d}$, $\operatorname*{\mathbb{P}}\{Y-\operatorname*{\mathbb{E}}[Y\mid\boldsymbol{X}=\boldsymbol{x}]>t\mid\boldsymbol{X}=\boldsymbol{x}\}\leq C\cdot\exp(-t^{2}/(2\sigma_{Y}^{2}))$.
@@
$d(\widehat{\boldsymbol{U}},\boldsymbol{U})\leq C_{1}\Bigg{[}{C_{h,\sigma_{\theta},1}\cdot{d(\mu_{f}+\sigma_{Y})^{2}}}\cdot\frac{\sqrt{m}}{n}+{C_{h,\sigma_{\theta},2}\cdot{d^{1/2}(\mu_{f}+\sigma_{Y})}}\cdot\frac{1}{\sqrt{n}}\ \Bigg{]}\cdot\frac{\log({8d}/{{\delta}})}{\sqrt{\delta}}\,,$ (6)
@@
$C_{h,\sigma_{\theta},2}$ $:=(1-10\sigma_{\theta}^{2}/h^{2})^{-d/2}\cdot\mu_{\nabla}\cdot\mu_{\rho}^{5/6}\cdot(\lambda_{k}(\overline{\boldsymbol{M}}))^{-1}.$ (7)
@@
The dependence on the failure probability $\delta$ is of order $\tilde{O}(\delta^{-1/2})$ which may appear large in light of usual high-probability results. This is due to the fact that we make minimal distributional assumptions, which in particular result in lower order moment conditions on the random matrices $\hat{\boldsymbol{\beta}}_{j,1}\hat{\boldsymbol{\beta}}_{j,2}^{{\scriptscriptstyle{\mathsf{T}}}}$, preventing exponential concentration guarantees. Nonetheless, heavy tail approaches such as median-of-means (see e.g. [17]) may be employed to improve such dependence on $\delta$.

### v2

@@ Abstract
We consider the problem of sufficient dimension reduction (SDR) for multi-index models. The estimators of the central mean subspace in prior works either have slow (non-parametric) convergence rates, or rely on stringent distributional conditions (e.g., elliptical symmetric covariate distribution $P_{\boldsymbol{X}}$). In this paper, we show that a fast parametric convergence rate of form $C_{d}\cdot n^{-1/2}$ is achievable via estimating the expected smoothed gradient outer product, for a general class of distribution $P_{\boldsymbol{X}}$ that admits Gaussian or heavier distributions. When the link function is a polynomial with a degree of at most $r$ and $P_{\boldsymbol{X}}$ is the standard Gaussian, we show that the prefactor depends on the ambient dimension $d$ as $C_{d}\propto d^{r}$.
@@ 1 Introduction
Our result relies on estimating outer products of a smoothed version of $\nabla f$, where the smoothing itself—induced by careful reweighting of the training sample—leverages a first-order Stein’s Lemma to guarantee full recovery of $\operatorname{col}(\boldsymbol{U})$. Similar to other existing analyses, we use the knowledge of $P_{\boldsymbol{X}}$ to properly reweight the training sample.
@@ 1.1 Prior Theoretical Works
As the name suggests, the class of inverse regression methods leverage functionals of the inverse relation $\boldsymbol{X}\mid Y$ that may reveal the index space or a subspace thereof. Although these methods are initially designed to estimate the central subspace, they can also serve the purpose of estimating the central mean subspace when the covariates $\boldsymbol{X}$ and the error $\varepsilon=Y-g(\boldsymbol{X})$ are independent, under which case these two subspaces are identical. For instance, Sliced Inverse Regression (SIR) by [23] estimates $\boldsymbol{M}=\text{Cov}[\operatorname*{\mathbb{E}}[\boldsymbol{X}\mid Y]]$ and Sliced Average Variance Estimator (SAVE) by [10] estimates $\boldsymbol{M}=\operatorname*{\mathbb{E}}[(\boldsymbol{I}_{d}-\text{Cov}[\boldsymbol{X}\mid Y])^{2}]$. By carefully estimating means or covariances over slices, i.e., partitions of the $Y$ space, parametric convergence rates of the form $O(n^{-1/2})$ may be guaranteed, albeit under restrictive distributional conditions.
For instance, SIR imposes a so-called linear conditional mean (LCM) condition (where $\operatorname*{\mathbb{E}}[\boldsymbol{X}\mid\boldsymbol{U}^{{\scriptscriptstyle{\mathsf{T}}}}\boldsymbol{X}]$ is linear in $\boldsymbol{U}^{{\scriptscriptstyle{\mathsf{T}}}}\boldsymbol{X}$) in order to ensure $\operatorname{col}(\boldsymbol{M})\subset\operatorname{col}(\boldsymbol{U})$, i.e., in order to guarantee partial recovery of $\operatorname{col}(\boldsymbol{U})$; if in addition we impose a so-called random conditional projection (RCP) condition (requiring for all $v\in\mathbb{R}^{d}$ that $\operatorname*{\mathbb{E}}[\boldsymbol{v}^{{\scriptscriptstyle{\mathsf{T}}}}\boldsymbol{X}\mid Y]$ is not constant as a r.v.) then one can ensure $\operatorname{col}(\boldsymbol{M})=\operatorname{col}(\boldsymbol{U})$. These distributional conditions, by restricting the relation between $P_{\boldsymbol{X}}$ and the index space $\operatorname{col}(\boldsymbol{U})$, rule out some rather benign settings: for instance, for $k=1$, it is well known that under Gaussian $P_{\boldsymbol{X}}$, these conditions cannot hold for even link functions such as $f(z)=z^{2}$ (also referred as phase retrieval problem). Similar examples of failure modes are discussed for SAVE in Appendix A. Additional discussions and examples for non-exhaustiveness issues can be found in [16] and [21].
@@
Table 1: Various Methods of Estimating CMS with Their Key Assumptions and Theoretical Guarantee.
Methods Single/Multiple Key Error Prefactor with
Index Assumptions Rate Polynomial Link
SIR[23] Multiple LCM, RCP $O(n^{-1/2})$ NA
SAVE[10] Multiple LCM, RCP $O(n^{-1/2})$ NA
MAVE[31] Multiple Smooth Link Functions $O(n^{-1/\Omega(k)})$ NA
SAA[17, 18] Multiple Smooth Link Functions $O(n^{-1/2})$ NA
for $k\leq 3$
SOSM[32] Multiple Non-degenerate Hessian $O(n^{-1/2})$ NA
ADE[4] Single Gaussian Covariates $O(n^{-1/2})$ NA
EIVH[16] Single Gaussian Covariates $O(n^{-1/2})$ NA
Smooth Link Function
ShallowNet[3] Single Gaussian Covariates $O(n^{-1/2})$ $C_{d}\sim d^{r/2}$
Smooth Link Function
SmoothedSGD[14] Single Gaussian Covariates $O(n^{-1/2})$ $C_{d}\sim d^{r/4}$
Known Link Function
ESGOP Multiple Known Density Ratio $O(n^{-1/2})$ $C_{d}\sim d^{r}$
(Our Method) Smooth Link Function
Finally, for easier reference, we list the key assumptions and theoretical guarantee of some popular works in Table 1.
@@ 1.2 Detailed Contributions
Our main result states that, for a general class of distributions $P_{\boldsymbol{X}}$ admitting Gaussians and heavier tail distributions, our estimator admits a parametric error rate of the form $C_{d}\cdot n^{-1/2}$. Finally, to elucidate the leading constant $C_{d}$, we consider the special case of polynomial links $g$—which has received much recent interest—and show that it behaves as $d^{O(r)}$ for polynomials of degree at most $r$ (Corollary 3.20), in line with recent lower-bounds of [13] for certain classes of computationally tractable procedures.
@@ 2.2 Central Mean Subspace
(1) $g(\boldsymbol{X})=f(\boldsymbol{U}^{{\scriptscriptstyle{\mathsf{T}}}}\boldsymbol{X})$
@@
Given an orthogonal matrix $\boldsymbol{U}$ and its estimate ^$\widehat{}\boldsymbol{U}$ (which is also orthogonal), the distance with optimal rotation between them defined as:
$d(\widehat{}\boldsymbol{U},\boldsymbol{U}):=\min_{\boldsymbol{R}\in\mathcal{O}^{k\times k}}\lVert\widehat{}\boldsymbol{U}\boldsymbol{R}-\boldsymbol{U}\rVert_{\operatorname{op}}\,,$
@@
$c\cdot d(\widehat{\boldsymbol{U}},\boldsymbol{U})\leq\sin_{\Theta}(\widehat{\boldsymbol{U}},\boldsymbol{U})\leq C\cdot d(\widehat{\boldsymbol{U}},\boldsymbol{U}).$
@@ 3 Main Results
We first introduce a surrogate object $\overline{\boldsymbol{M}}$ in Section 3.1 below, and show that it indeed is exhaustive, i.e., it recovers all of $\operatorname{col}({\bf U})$. Following this, we derive a separate surrogate object $\widetilde{\boldsymbol{M}}$ in Section 3.3 which approximates $\overline{\boldsymbol{M}}$ and turns out to be simpler to estimate and retains the same exhaustiveness properties. Our main algorithm then estimates $\widetilde{\boldsymbol{M}}$, and returns the column space ^$\hat{}\boldsymbol{U}$ of this estimate. Our main result of Theorem 3.14 which bounds the estimation error is presented in Section 3.4, followed by an instantiation for polynomial link functions in Section 3.5.
@@ 3.1 Expected Smoothed Gradient Outer Product
(2) $\overline{\boldsymbol{M}}:=\operatorname*{\mathbb{E}}_{\boldsymbol{\theta}\sim\mathcal{N}(\boldsymbol{0}_{d},\sigma_{\theta}^{2}\boldsymbol{I}_{d})}\left[\boldsymbol{\beta}_{h}(\boldsymbol{\theta})\boldsymbol{\beta}_{h}(\boldsymbol{\theta})^{{\scriptscriptstyle{\mathsf{T}}}}\right]=\boldsymbol{U}\left(\operatorname*{\mathbb{E}}_{\boldsymbol{\theta}\sim\mathcal{N}(\boldsymbol{0}_{d},\sigma_{\theta}^{2}\boldsymbol{I}_{d})}\left[\overline{\nabla}_{h}f(\boldsymbol{\theta})\overline{\nabla}_{h}f(\boldsymbol{\theta})^{\scriptscriptstyle{\mathsf{T}}}\right]\right)\boldsymbol{U}^{\scriptscriptstyle{\mathsf{T}}}.$
@@ 3.2 Average Smoothed Gradient Outer Product
(3) $\widetilde{\boldsymbol{M}}:=\frac{1}{m}\sum_{j=1}^{m}\boldsymbol{\beta}_{h}(\boldsymbol{\theta}_{j})\boldsymbol{\beta}_{h}(\boldsymbol{\theta}_{j})^{{\scriptscriptstyle{\mathsf{T}}}}=\boldsymbol{U}\left(\frac{1}{m}\sum_{j=1}^{m}\left[\overline{\nabla}_{h}f(\boldsymbol{\theta}_{j})\overline{\nabla}_{h}f(\boldsymbol{\theta}_{j})^{\scriptscriptstyle{\mathsf{T}}}\right]\right)\boldsymbol{U}^{\scriptscriptstyle{\mathsf{T}}}.$
@@
The smoothed gradient ¯$\overline{\nabla}_{h}f$ satisfies:
@@ 3.3 Estimation of CMS
For intuition, notice that we have $\boldsymbol{\beta}_{h}(\boldsymbol{\theta})=\operatorname*{\mathbb{E}}_{\boldsymbol{Z}\sim\mathcal{N}(\boldsymbol{0},h^{2}\boldsymbol{I}_{d})}\nabla g(\boldsymbol{Z}+\boldsymbol{\theta})=\operatorname*{\mathbb{E}}_{\boldsymbol{Z}\sim\mathcal{N}(\boldsymbol{\theta},h^{2}\boldsymbol{I}_{d})}\nabla g(\boldsymbol{Z})$.
Thus, suppose for a moment we were to sample $\boldsymbol{X}\sim\mathcal{N}(\boldsymbol{\theta},h^{2}\boldsymbol{I}_{d})$, then we would have
@@
(4) $\rho_{h}(\boldsymbol{x};{\boldsymbol{\theta}}):=\frac{\varphi_{h}(\boldsymbol{x}-\boldsymbol{\theta})}{p(\boldsymbol{x})}\,,\quad\text{where }\varphi_{h}(\cdot)\text{ denotes the density of $\mathcal{N}(\boldsymbol{0}_{d},h^{2}\boldsymbol{I}_{d})$ }.$
@@
(5) $\hat{\boldsymbol{\beta}}_{h}(\boldsymbol{\theta}_{j};\mathcal{D}_{j,\ell}):=\frac{h^{-2}}{|\mathcal{D}_{j,\ell}|}\sum_{(\boldsymbol{X},Y)\in\mathcal{D}_{j,\ell}}\rho_{h}(\boldsymbol{X};\boldsymbol{\theta}_{j})\cdot Y\cdot(\boldsymbol{X}-\boldsymbol{\theta}_{j}).$
@@
The estimators defined in Equation 5 depend on the knowledge of the density ratio $\rho_{h}$. For the cases of unknown $\rho_{h}$, one may consider a plug-in estimator where $\rho_{h}$ is replaced by its estimates $\hat{\rho}_{h}$ that is calculated with another unlabeled dataset $\{\widetilde{\boldsymbol{X}}_{1},\ldots,\widetilde{\boldsymbol{X}}_{N}\}$. Such a strategy works best when one has access to abundant unlabeled data while the budget for labeling is limited. For a more detailed discussion of the effect of not knowing $\rho_{h}$, please refer to Appendix D.
Remark 3.9.
@@
Notice that, by the independence of the two estimates of $\boldsymbol{\beta}_{h}(\boldsymbol{\theta}_{j})$, we also immediately have that $\operatorname*{\mathbb{E}}[\hat{\boldsymbol{\beta}}_{j,1}\hat{\boldsymbol{\beta}}_{j,2}^{{\scriptscriptstyle{\mathsf{T}}}}\mid\boldsymbol{\theta}_{j}]={\boldsymbol{\beta}}_{h}(\boldsymbol{\theta}_{j}){\boldsymbol{\beta}}_{h}(\boldsymbol{\theta}_{j})^{{\scriptscriptstyle{\mathsf{T}}}}$. This leads to the estimator of $\widetilde{\boldsymbol{M}}$ given below:
@@ 3.4 Upper-bound on Estimation Error
The noise has a subgaussian tail, i.e., $\exists C,\sigma_{Y}^{2}>0$ such that $\forall\boldsymbol{x}\in\mathbb{R}^{d}$, $\operatorname*{\mathbb{P}}\{|Y-\operatorname*{\mathbb{E}}[Y\mid\boldsymbol{X}=\boldsymbol{x}]|>t\mid\boldsymbol{X}=\boldsymbol{x}\}\leq C\cdot\exp(-t^{2}/(2\sigma_{Y}^{2}))$.
@@
(6) $d(\widehat{\boldsymbol{U}},\boldsymbol{U})\leq C_{1}\Bigg{[}{C_{h,\sigma_{\theta},1}\cdot{d(\mu_{f}+\sigma_{Y})^{2}}}\cdot\frac{\sqrt{m}}{n}+{C_{h,\sigma_{\theta},2}\cdot{d^{1/2}(\mu_{f}+\sigma_{Y})}}\cdot\frac{1}{\sqrt{n}}\ \Bigg{]}\cdot\frac{\log({8d}/{{\delta}})}{\sqrt{\delta}}\,,$
@@
(7) $C_{h,\sigma_{\theta},2}$ $:=(1-10\sigma_{\theta}^{2}/h^{2})^{-d/2}\cdot\mu_{\nabla}\cdot\mu_{\rho}^{5/6}\cdot(\lambda_{k}(\overline{\boldsymbol{M}}))^{-1}.$
@@
The dependence on the failure probability $\delta$ is of order $\tilde{O}(\delta^{-1/2})$ which may appear large in light of usual high-probability results. This is due to the fact that we make minimal distributional assumptions, which in particular result in lower order moment conditions on the random matrices $\hat{\boldsymbol{\beta}}_{j,1}\hat{\boldsymbol{\beta}}_{j,2}^{{\scriptscriptstyle{\mathsf{T}}}}$, preventing exponential concentration guarantees. Nonetheless, heavy tail approaches such as median-of-means (see Appendix E) may be employed to improve such dependence on $\delta$.